OpenStudy (anonymous):

Can anyone explain this to me? I have no idea how to find this answer... 6a+8-14a=96 Using Solving Multi-Step Equations

OpenStudy (anonymous):

You simply isolate the variable a by sending over the constants.

OpenStudy (anonymous):

Firstly, you regroup similar terms, in this case, the as.

OpenStudy (anonymous):

-8a+8=96

OpenStudy (anonymous):

-8a=96-8 -8a=88 a=88/-8 a=-11

OpenStudy (anonymous):

Thanks